# Contextuality of the probability current in quantum mechanics

### v1

@@ 3 Conclusion
We therefore have a situation where, while the various Lorentz observers agree on the values of the probabilities of all measurement results, they may disagree on the path taken by the probability to reach the detectors. Curiously, while they agree when they calculate the flux of the probability current (more precisely, its integral over time), they disagree on the lines of current that lead to this flux. The major difference of points of view occurs between observers with positive velocity $v$ and those with negative velocity. When $v$ vanishes, a discontinuous change occurs, which is reminiscent of the discontinuity taking place in Galilean relativity when the experimental setup is changed. Nevertheless, a discontinuity of the description of the same experiment by different observers raises a more serious problem, since it cannot be solved by invoking the Bohrian wholeness of the experimental apparatus. It therefore seems difficult to attribute physical reality to the probability current, and consequently to the probability fluid itself.
@@
All our reasoning took place within standard quantum mechanics, but of course it has consequences if one uses the dBB theory. A pointed out in Ref. [6], the so called quantum equilibrium condition of this theory cannot be relativistically invariant. One possibility is to decide that the theory requires the choice of a privileged frame, for instance that of the laboratory (but, even with this rule, problems arise when one considers experimental setups containing rapidly moving components [13, 14]) or the frame of the cosmic microwave radiation.
@@
But one can also take another point of view. The dBB theory achieves two main results: first, it unifies the dynamics, considering that measurements are just ordinary quantum interaction processes, and eliminating the need for any specific postulate (von Neumann projection, Heisenberg cut, etc.); second, it provides a description of reality (an ontology) expressed in terms of point positions of particles. Clearly, the second result is incompatible with a completely relativistic theory [^1] In the context of field theory, Bell [15] discusses why local occupation numbers (beables) for a fermionic field cannot provide a fully relativistic view of reality either., where no Lorentz frame is privileged. But, even if one gives up the description of reality with point particles, one can still preserve the first result: the positions are then considered just as an indicator pointing to the branch of the state vector that is active, and eliminates all the empty waves. In this way, the unification of the dynamics is preserved, and the closest description of reality is obtained in terms of the active branch of the state vector. True, the indicator that determines this active branch is not relativistically invariant but, since it appears as a purely mathematical variable in the dynamics, this does not necessarily ruin the approach.
@@
In conclusion, and whether or not a unified dynamics is desirable in quantum mechanics, it is clear that even blurred relativistic trajectories seem to be incompatible with this theory – this remains true if they are blurred more than what is requested by the Heisenberg uncertainty relations. Bohr went even further [16] by claiming that, due to the non-zero value of the Planck constant, “the space-time description and the claims of causality are of complementary nature”, and therefore incompatible; this should indeed forbid a description of the motion of the probability fluid in space-time. Discontinuities in the trajectories taking place when the Lorentz reference frame is changed may be seen as a particular case of application of this general statement.

### v2

@@ 3 Conclusion
We therefore have a situation where, while the various Lorentz observers agree on the values of the probabilities of all measurement results, they may disagree on the path taken by the probability to reach the detectors: different motions of the probability fluid in their respective frames are predicted by the theory. Technically, while they agree when they calculate the flux of the probability current (more precisely, its integral over time), they disagree on the lines of current that lead to this flux. The major difference of points of view occurs between observers with positive velocity $v$ and those with negative velocity. When $v$ vanishes, a discontinuous change occurs, which is reminiscent of the discontinuity taking place in Galilean relativity when the experimental setup is changed. Nevertheless, a discontinuity of the description of the same experiment by different observers raises a more serious problem, since it cannot be solved by invoking the Bohrian wholeness of the experimental apparatus. It therefore seems difficult to attribute physical reality to the probability current, and consequently to the probability fluid itself. In other words, the theory tells us where and when the probability arrives, but not how it propagates in space, which is not compatible with the motion of a fluid.
@@
All our reasoning took place within standard quantum mechanics, but of course it has consequences if one uses the dBB theory. As pointed out in Ref. [7], the so called quantum equilibrium condition of this theory cannot be relativistically invariant; see also Refs. [13, 5]. One possibility is to decide that the theory requires the choice of a privileged frame, for instance that of the laboratory (but, even with this rule, problems arise when one considers experimental setups containing rapidly moving components [15, 16]) or the frame of the cosmic microwave radiation.
@@
One may consider even less orthodox interpretations of quantum mechanics, and take another point of view emphasizing waves instead of point positions. The dBB theory achieves two main results: first, it unifies the dynamics, considering that measurements are just ordinary quantum interaction processes, and eliminating the need for any specific postulate (von Neumann projection, Heisenberg cut, etc.); second, it provides a description of reality (an ontology) expressed in terms of point positions of particles. Clearly, the second result is incompatible with a completely relativistic theory [^4] In the context of field theory, Bell [17] discusses why local occupation numbers (beables) for a fermionic field cannot provide a fully relativistic view of reality either., where no Lorentz frame is privileged. But, even if one gives up the description of reality with point particles, one can still preserve the first result: the positions are then not considered as real, but just as a mathematical indicator pointing to the branch of the state vector that is active, eliminating the empty waves. In this unorthodox “reverse Bohmian interpretation”, the unification of the dynamics is preserved, and the closest description of reality is obtained in terms of the active branch of the state vector. This view is reminiscent of Schrödinger’s initial hopes, considering the waves as real, while it requires to accept that these waves propagate in configuration space (instead of ordinary space). It could also be seen as a variant of the Everett interpretation, but without many parallel versions of the world; this simplification eliminates the difficult definition of the probabilities of human perceptions in the many worlds, and may make the theory more acceptable. True, the indicator that determines this active branch is then not relativistically invariant; but, since it appears as a purely mathematical variable in the dynamics, this does not necessarily ruin the approach.
@@
In conclusion, and whether or not a unified dynamics is desirable in quantum mechanics, it is clear that even blurred relativistic trajectories seem to be incompatible with this theory – this remains true if they are blurred much more than what is requested by the Heisenberg uncertainty relations. Bohr went even further [18] by claiming that, due to the non-zero value of the Planck constant, “the space-time description and the claims of causality are of complementary nature”, and therefore incompatible; this should indeed forbid a description of the motion of the probability fluid in space-time. Discontinuities in the trajectories taking place when the Lorentz reference frame is changed may be seen as a particular case of application of this general statement.